# SGPMIL: Sparse Gaussian Process Multiple Instance Learning

### v1

@@ 5.4 Inducing Points Capture Patch Variability
Recent MIL work has highlighted the importance of bag-level representations, noting that noisy instance features can negatively impact classification performance [46]. In our approach, the learnable inducing points act as task-adaptive prototypes that capture salient morphological patterns and core structural features across tissue slides, as shown in Figure 3.
In Figure 3(a), we visualize the slide-level ground truth annotations. In (b), we show a label map where each patch embedding is assigned the label of its most similar inducing point (based on cosine similarity). This visualization reveals that distinct inducing points specialize in capturing different tissue types—e.g., tumor regions, dense stromal tissue, and interface zones such as tumor-stroma boundaries. Finally, in Figure 3(c), we display the top-7 most similar patches for the inducing points that match the largest number of patches. The visual coherence across patches associated with each inducing point demonstrates that the model learns meaningful and morphologically consistent prototypes.
@@ 5.5 Prediction Uncertainty Estimation
Figure 4: Predictive uncertainty for correctly and incorrectly classified WSIs in the BRACS (left) and PANDA (right) datasets. Each pair of boxplots represents the distribution of standard errors of the predicted class probabilities for correct (green) and incorrect (red) predictions. A statistically significant difference in uncertainty is observed for both datasets ($p<0.05$), based on a two-sample Welch’s $t$-test [40].
In Figure 4, we visualize the prediction uncertainty for correctly and incorrectly classified WSIs. Specifically, we compute the standard deviation of the predicted class probabilities across multiple samples and compare the resulting distributions using a two-sample Welch’s $t$-test [40]. We consider BRACS and PANDA datasets as for CAMELYON16 and TCGA-NSCLC, the incorrectly predicted slides are very few to properly estimate and compare the variance between sets (high ACC, AUC). We find a statistically significant difference ($p<0.05$) in predictive uncertainty between correct and incorrect predictions across datasets. This indicates that higher uncertainty is associated with misclassified slides, suggesting that our model’s uncertainty estimates are well-calibrated—an essential property for deployment in safety-critical clinical workflows.
@@ 6 Conclusion
We introduced SGPMIL, a robust and scalable probabilistic attention-based MIL framework built on sparse Gaussian processes, designed for improved instance- and bag-level prediction in high-dimensional histopathology tasks. By learning a variational posterior over attention scores and leveraging a flexible kernel, SGPMIL captures semantic relationships among instances while enabling principled uncertainty quantification. Our method achieves leading instance-level performance, competitive bag-level accuracy, and addresses key limitations of prior probabilistic MIL approaches, particularly in terms of numerical stability and scalability. Our code will be made publicly available to support reproducibility and further research. A limitation of this work is that we focus exclusively on unimodal visual inputs. Extending SGPMIL to multimodal MIL settings—such as integrating clinical or genomic metadata alongside WSIs—remains an exciting direction for future research.
Acknowledgements
This work has been partially supported by project MIS 5154714 of the National Recovery and Resilience Plan Greece 2.0 funded by the European Union under the NextGenerationEU Program. Hardware resources were granted with the support of GRNET.

### v2

@@ 5.4 Inducing Points Capture Patch Variability
Recent MIL work has highlighted the importance of bag-level representations, noting that noisy instance features can negatively impact classification performance [60]. In our approach, the learnable inducing points act as task-adaptive prototypes that capture salient morphological patterns and core structural features across tissue slides, shown in Fig. 3.
In Figure 3(a), we visualize the slide-level ground truth annotations. In (b), we show a label map where each patch embedding is assigned the label of its most similar inducing point (based on cosine similarity). This visualization reveals that distinct inducing points specialize in capturing different tissue types—e.g., tumor regions, dense stromal tissue, and interface zones such as tumor-stroma boundaries. Finally, in Figure 3(c), we display the top-7 most similar patches for the inducing points that match the largest number of patches. The visual coherence across patches associated with each inducing point demonstrates that the model learns meaningful and morphologically consistent prototypes. Additional inducing point label maps and HoverNet-based cancer/nucleus-type distributions [18] are shown in Supplementary Figures 4, 5 and 6.
@@ 5.5 Prediction Uncertainty Estimation
Figure 4: Prediction uncertainty for correctly (green) and incorrectly (red) classified WSIs in the BRACS (left) and PANDA (right) datasets. Each boxplot pair shows the distribution of standard deviation of predicted class probabilities. A statistically significant difference in uncertainty is observed in both datasets ($p<0.05$, Welch’s $t$-test [53]).
In Figure 4, we visualize the prediction uncertainty for correctly and incorrectly classified WSIs. We compute the standard deviation of the predicted class probabilities across samples and compare the resulting distributions using a two-sample Welch’s $t$-test [53]. We consider BRACS and PANDA datasets as CAMELYON16 and TCGA-NSCLC contain too few misclassified slides for reliable variance estimation (due to high ACC and AUC). A statistically significant difference ($p<0.05$) in prediction uncertainty is observed between correct and incorrect predictions across datasets. This indicates that higher uncertainty is associated with misclassified slides, suggesting that our model’s uncertainty estimates are well-calibrated—an essential property for deployment in safety-critical clinical workflows.
@@ 6 Conclusion
We introduced SGPMIL, a robust and scalable probabilistic attention-based MIL framework built on SGP, designed for improved instance- and bag-level prediction in histopathology tasks. By learning a variational posterior over attention scores and leveraging a flexible kernel, SGPMIL captures semantic relationships among instances while enabling principled uncertainty quantification. Our method achieves leading instance-level performance, competitive bag-level accuracy, and addresses key limitations of prior probabilistic MIL approaches, particularly in terms of numerical stability and scalability. A limitation of this work is that we focus exclusively on unimodal visual inputs. Extending SGPMIL to multimodal MIL settings—such as integrating clinical or genomic metadata alongside WSIs—remains an exciting direction for future research.
Acknowledgments
This work has been partially supported by project MIS 5154714 of the National Recovery and Resilience Plan Greece 2.0 funded by the European Union under the NextGenerationEU Program. Hardware resources were granted with the support of GRNET. We also gratefully acknowledge the DATAIA program for supporting JD as a visiting professor at Université Paris-Saclay.